40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1649     Entry time: Sat Jun 25 18:12:56 2016
Author: Antonio 
Type: DailyProgress 
Category: optic 
Subject: North path construction progress 

North Path

SUMMARY

The north path is going to be redesign because we want to be able to implement the PMC once we have all the electronics ready.

There is no room for an additional EOM since one more EOM is required to produce sidebands to lock the PMC with the PDH method.

Thigs done (Following the attached schematic):

  1. The lens (77.3, f = 171.9mm) before the FI has been removed and placed after it;
  2. The FI is closer and a bit tilted as the the back reflection was going straight to the laser;
  3. lambda/4 and lambda/2 have been placed in the path before the FI and have been rotated in a way that at the output of the FI we have maximum power (~0.7W);
  4. A lens (77.3) has been placed after the FI;
  5. Lambda/2 has been placed in order to create P polarized light;
  6. A high power PBS (PBSO borrowed from the 40m) has been implemented for power adjust;
  7. Lambda/2 after the PBSO in order to create s-polarized light;
  8. Beam has been measured in the area after the lambda/2 (distance reference) (as shown in the picture);
  9. A BroadBand (BBEOM) has been placed with waist in the middle with resulting beam aligned in plus menus 10 micro meter range (This means that looking at a reference point with the ccd camera, after BBEOM was placed and aligned the beam was hitting the same point. ) The power of the incoming beam was 320uW while the power of the outcoming beam was 310uW.
  10.  A lens (51.5, f = 114.5mm) after the BBEOM has been cleaned and placed (This lens has a little scratch);
  11. A steering mirror (s) has been cleaned and implemented;
  12.  A lambda/2 has placed after the mirror in order to have p-light;
  13.  The AEOM has been implemented;
  14.  Lambda/4 has been implemented and rotated at 10deg in order to have circular polarized light;
  15.  PBS has been cleaned and implemented;
  16.  Lambda/2 has been implemented in order to give s-polarized light (45deg);
  17.  A lens (51.5, f = 114.5mm) has been cleaned and implemented in order to create the right beam for the resonant EOM to be used for the PMC sidebands;
  18. Beam profile has been measured: the fit is not ok but the data are indicative enough (waist of ~ 100 um);
  19.  A steering mirror (not labeled) has been power checked (“equal” for both s and p) and implemented;
  20.  The resonant EOM (21.5MHz) to be used for the PMC sidebands has been implemented with alignment precision at the order of tens of microns (incoming power was 315 uW while the outcoming power was 306 uW);
  21. Steering mirror implemented;
  22.  Currently we are in the process of implementing two lenses to mode match the PMC;

Things to note:

  1. The beam coming out of the laser is not perfectly at 3 inches height;
  2. The resonant EOM used for the PMC will be removed while we take measurement without PMC;
  3.  The beam coming out of the laser is very dirty;
  4. There is a specific lens mount which fails to have the optics center at 1 inch;

Next step:

  1. Mode matching for the PMC;

Attachment:

1. Schematic 

     2. Beam profile after l/2 (see point 8, zero is at l/2)

     

   3. Beam profile measured before EAOM (zero at the steering mirror)

4. Beam profile measure before EOM for PMC (zero reference at lens (51.5))

(to be updated: the fit is failing)

Attachment 1: North_layout_progress.graffle  27 kB
Attachment 2: North_layout_progress.graffle  27 kB
Attachment 3: North_layout_progress.pdf  64 kB  | Hide | Hide all
North_layout_progress.pdf
Attachment 4: beam_profile_after_lambda_2.pdf  14 kB  | Hide | Hide all
beam_profile_after_lambda_2.pdf
Attachment 5: beam_profile_before_AEOM.pdf  13 kB  | Hide | Hide all
beam_profile_before_AEOM.pdf
ELOG V3.1.3-